# Littlewood-Offord bounds on the symmetric groups and applications

### v1

@@ 2.4. Paper Organization
Let $S_{\pi}((b_{ij}))=\sum_{i}b_{i\pi(i)}$, where $\pi$ is a uniform permutation. We have ${\mathbb{E}}S_{\pi}((b_{ij}))=(1/n)\sum_{ij}b_{ij}$. It can also be shown that [16, formula (89)]
$\operatorname{Var}S_{\pi}((b_{ij}))=[1/4n^{2}(n-1)]\sum_{i\neq j;k\neq l}(b_{ik}-b_{jk}-b_{il}+b_{jl})^{2}.$ (18)
So if we assume that $b_{ik}-b_{jk}-b_{il}+b_{jl}\in[-N,N]$ then
$\operatorname{Var}(S_{\pi}((b_{ij})))\leq nN^{2}.$
It follows by Chebyshev that with probability at least $1-16/C^{2}$,
$|S_{\pi}((b_{ij}))-{\mathbb{E}}S_{\pi}((b_{ij}))|\in[-(C/4)\sqrt{n}N,(C/4)\sqrt{n}N].$ (19)
For each fixed $1\leq s\leq r$, apply this bound to the sequence $b_{ij}=a_{ij;s}$ from (17), and take the intersection, we see that there is an event $\mathcal{E}$} (over the random permutation $\pi$) with ${\mathbb{P}}({\mathcal{E}})\geq 1-16r/C^{2}$ on which, simultaneously for all $s$,
$|S_{\pi}((a_{ij;s}))-\mu_{s}|\in[-C\sqrt{n}N_{s},C\sqrt{n}N_{s}],$
where $\mu_{s}={\mathbb{E}}S_{\pi}((a_{ij;s}))$. Under this event $\mathcal{E}$}, the original sum $\sum_{i}a_{i\pi(i)}^{\prime}$ belongs to the shifted GAP
$\sum_{i}a_{i\pi(i)}^{\prime}\in\Big([\mu_{1}-C\sqrt{n}N_{1},\mu_{1}+C\sqrt{n}N_{1}]\cap{\mathbb{Z}}\Big)g_{1}+\dots+\Big([\mu_{r}-C\sqrt{n}N_{r},\mu_{r}+C\sqrt{n}N_{r}]\cap{\mathbb{Z}}\Big)g_{r}.$
By pigeonhole principle, there exists $x$ from this GAP such that
${\mathbb{P}}(S_{\pi}(a_{ij}^{\prime})=x)\geq\frac{{\mathbb{P}}({\mathcal{E}})}{(3C)^{r}n^{r/2}\prod N_{i}}\geq\rho,$
with appropriate choices of the constants. ∎
@@
Embedding. The following theorem (see [56, Lemma 5.25], [39, Theorem 4.3]) allows us to assume that $a_{ij}$ are elements of ${\mathbb{F}}_{p}$ for some large prime $p$.
@@
Let $n^{\prime}={\varepsilon}|A|={\varepsilon}n^{4}$. By averaging, at least a $(1-{\varepsilon})$-fraction of $a\in A$ satisfy
$\sum_{t\in L_{m}}\|\frac{at}{p}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\leq\frac{4mn^{3}}{n^{\prime}}|L_{m}|.$
@@
$\sum_{t\in L_{m}}\|\frac{at}{p}\|^{2}_{{\mathbb{R}}/{\mathbb{Z}}}\leq l^{2}\frac{4mn^{3}}{n^{\prime}}|L_{m}|.$ (20)
@@
Set $k=\sqrt{\frac{n^{\prime}}{160mn^{3}}}$. By (20), we have $\bigcup_{l=1}^{k}lA^{\prime}\subset L_{m}^{*}$. Setting $A^{\prime\prime}=A^{\prime}\cup\{0\}$, this implies $kA^{\prime\prime}\subset L_{m}^{*}\cup\{0\}$. Hence
@@
The ambient field ${\mathbb{F}}_{p}$ is no longer important, so we may view the $a_{ij}$ as integers. We now invoke the following long-range inverse theorem (see [39, Theorem 3.2]).
@@
$k=\Omega\Big(\sqrt{\frac{n}{m}}\Big)=\Omega\Big(\sqrt{\frac{n}{\log n}}\Big),\quad\rho^{-1}\leq n^{C}\leq k^{2C+1},$
(22) allows us to apply Theorem 3.4 with $\gamma=2C+1$ and $X$ the set of distinct elements of $A^{{}^{\prime\prime}}$ (note that $kX=kA^{{}^{\prime\prime}}$ for $k\geq 2$). Hence $X$ is contained in a proper symmetric GAP $Q$ of rank $r=O_{C,{\varepsilon}}(1)$ and size
$O_{C,{\varepsilon}}(k^{-r}|kX|)=O_{C,{\varepsilon}}(k^{-r}|kA^{{}^{\prime\prime}}|)=O_{C,{\varepsilon}}\Big(\rho^{-1}e^{-m+2}(\sqrt{\frac{n}{m}})^{-r}\Big)=O_{C,{\varepsilon}}(\rho^{-1}n^{-r/2}),$
@@
To complete the section we quickly deduce Theorem 1.23 and Theorem 1.24.
@@
(of Theorem 1.23) Assume for contradiction that $\rho:=\sup_{x}{\mathbb{P}}\big(\sum_{i=1}^{n}w_{i}v_{\pi(i)}=x\big)\geq An^{-3/2}$ for some sufficiently large constant $A$. Set ${\varepsilon}=.1$ and $C=3$. By Theorem 1.11, there is a GAP $Q$ of rank $r$ and size $O(\rho^{-1}n^{-r/2})$ that contains a $(1-{\varepsilon})$-portion of the set of quadruples $\{(w_{i}-w_{j})(v_{k}-v_{l}):1\leq i,j,k,l\leq n\}$. This set must have at least $\Theta(n)$ (distinct) elements, because by assumption there are $\Theta_{c}(n^{2})$ pairs $i,j$ where $w_{i}-w_{j}\neq 0$, and on average, each such pair $(i,j)$ is associated with $\Theta(n^{2})$ pairs $(k,l)$. Among these pairs, we just chose $\Theta(n)$ pairs of the form $(k_{0},l_{1}),\dots,(k_{0},l_{m}),m=\Theta(n)$. Since $Q$ has rank at least $1$, we have $n\ll|Q|=O(\rho^{-1}n^{-1/2})=O(n/A)$, a contradiction if $A$ is sufficiently large. ∎
@@
(of Theorem 1.24) Let $S$ be the set $\{(w_{i}-w_{j})(v_{k}-v_{l})\}$ over a $(1-{\varepsilon})$-portion of the quadruples. We claim that $S$ must have at least $\Theta(n^{2}/\log n)$ elements. Since $Q$ has rank at least $1$, this would imply $n^{2}\log n\ll|Q|=O(\rho^{-1}n^{-1/2})$, which yields $\rho\ll n^{-5/2}\log n$, as desired.
It remains to prove $|S|\gg n^{2}/\log n$. Let $A=\{v_{1},\ldots,v_{n},w_{1},\ldots,w_{n}\}$. For $x\in A-A$, let $r(x)$ denote the number of representations ab=xa-b=x with $(a,b)\in A^{2}$. By [5, Theorem 3], [^2] The statement in [5] is given for $A\subset{\mathbb{Z}}$, but the proof works verbatim for $A\subset{\mathbb{R}}$. we have
$|A|^{6}\log|A|\gg\sum_{xy=x^{\prime}y^{\prime}}r(x)r(x^{\prime})r(y)r(y^{\prime})=\sum_{z}\Big|\sum_{xy=z}r(x)r(y)\Big|^{2}.$
@@
$|A|^{6}\log|A|\geq\frac{1}{|S|}\Big(\sum_{xy\in S}r(x)r(y)\Big)^{2}.$
Moreover, $\sum_{xy\in S}r(x)r(y)$ counts at least the number of selected quadruples, namely $(1-{\varepsilon})n^{4}$. Therefore, $|S|\gg n^{8}/|A|^{6}\log|A|=\Theta(n^{2}/\log n)$, as claimed. ∎
4. Continuous setting: proof of Theorems 1.15 and 1.17
@@
Fact 4.1.
Let $I\subset[n]$ with $|I|\geq\delta n$ for some constant $\delta>0$. Then, there exists ${\mathcal{R}}\subset\{-n,\dots,n\}$ of size $|{\mathcal{R}}|\gg_{\delta}n$ such that for every $r\in{\mathcal{R}}$, there are $\gg_{\delta}n$ pairs $x,y\in I$ with xy=rx-y=r.
Throughout this section, let $\mathcal{R}\subset\{-n,\dots,n\}$ be a set of size $|\mathcal{R}|=\Theta_{\delta}(n)$, as defined in Fact 4.1. Let $C$ and $A$ be positive constants, with $C$ sufficiently large relative to $\delta$ and $\varepsilon$, and $A$ sufficiently large relative to $C$.
@@
${\mathbb{P}}(|X|\leq 1)\ll\Big|\int_{|t|\leq 1}k(t)\varphi_{X}(t)\,dt\Big|\ll\int_{|t|\leq 1}|\varphi_{X}(t)|\,dt,$ (23)
@@
${\mathbb{P}}(|X|\leq 1)\ll\int_{|t|\leq 1}|\varphi_{S}(t)|\,dt\ll\int_{|t|\leq 1}\exp\Big\{-\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}dt.$ (24)
@@
$\sum_{r\in{\mathcal{R}}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\sum_{r\in{\mathcal{R}}}|t(w_{i}-w_{j})r|^{2}\asymp_{\delta}t^{2}n^{3}D_{k}^{2}.$
@@
$\sum_{r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}n.$
@@
$\sum_{(i,j)\in{\mathcal{G}},r\in{\mathcal{R}}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta,C}\sum_{k:\,|t|D_{k}\leq 1/(Cn)}t^{2}n^{3}D_{k}^{2}|{\mathcal{G}}_{k}|+\sum_{k:\,|t|D_{k}>1/(Cn)}n|{\mathcal{G}}_{k}|.$
@@
$\sum_{i,j,r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\gg_{\delta,C}An^{2}\log n.$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}(t(w_{i}-w_{j})r)^{2}\asymp_{\delta}t^{2}n^{2}.$
@@
$\int_{|t|\leq(\sqrt{A\log n})/n}|\varphi_{S}(t)|\,dt\leq\int_{|t|\leq(\sqrt{A\log n})/n}\exp(-\Theta(t^{2}n^{2}))\,dt\ll\frac{1}{n}\int_{{\mathbb{R}}}\exp(-\Theta(x^{2}))\,dx\ll\frac{1}{n}.$
@@
${\mathbb{P}}\Big(\Big|\sum_{i}w_{i}v_{\pi(i)}-L\Big|\leq\frac{1}{n}\Big)\ll\max\Big\{n^{-\sqrt{A}},\ \frac{1}{n}e^{-\Theta(L^{2})}\Big\}.$ (26)
@@
$\sum_{r\in{\mathcal{R}}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta}n.$
@@
$\sum_{i,j,r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\gg_{\delta}n|{\mathcal{G}}|\gg_{\delta}A^{2}n^{2}\log n.$
@@
Our plan in the proof of (6) is to replace $k(t)$ in (23) by some smoother function that can be extended holomorphically to $\mathbb{C}$}. Our starting point is that [^3] This identity is also frequently used to estimate small ball probabilities, see [56, Section 7].
@@
$\int_{(\sqrt{A\log n})/n\leq|t|\leq\sqrt{A\log n}}e^{-\pi t^{2}}\varphi_{X}(t)dt\leq(2\sqrt{A\log n})n^{-2\sqrt{A}}\ll n^{-\sqrt{A}}.$
@@
$\Big|e^{-\pi(t-icL)^{2}/n^{2}}\Big|=e^{-\pi(t^{2}-c^{2}L^{2})/n^{2}}\asymp 1,\quad\text{as }n\to\infty\text{ and }|t|\leq L=O(\sqrt{\log n}),$
@@
${\mathbb{P}}(|X|\leq 1)\leq\frac{e^{-cL^{2}}}{n}\cdot e^{C_{0}^{\prime}c^{2}L^{2}}\cdot 2L\ll\frac{e^{-L^{2}/(16C_{0}^{\prime})}\,L}{n}\ll\frac{e^{-\Theta(L^{2})}}{n}.$
@@
We observe that the assumption $|w_{i}-w_{j}|\leq\frac{1}{A\sqrt{\log n}}$ in Theorem 1.15 cannot be relaxed to $|w_{i}-w_{j}|\gg\frac{1}{\sqrt{\log n}}$ when using only the characteristic function method (i.e., relying solely on Theorem 2.1). Indeed  [^4] In this example, the matrix is very sparse, making the permanent-based bound weak. This suggests that a more refined approach may be needed, possibly restricting attention to the support of the $w_{i}$., suppose we partition $[n]$ into two disjoint sets $I\cup J=[n]$ with $|I|=2c\log n$ for some constant $c>0$. Let $w_{i}=\frac{1}{\sqrt{2c\log n}}$ and $\frac{1}{\sqrt{2c\log n}}$ for half of the $i\in I$ respectively, and $w_{j}=0$ for all $j\in J$. Then,
@@
$\sup_{x}{\mathbb{P}}(|S^{\prime}-x|\leq 1)\ll\int_{|t|\leq 1}|\varphi_{S^{\prime}}(t)|\,dt\ll\int_{|t|\leq 1}\exp\Big\{-\frac{1}{2n^{3}}\sum_{i,j,k,l}\Big\|\frac{tn^{3/2}}{\Delta}(w_{i}-w_{j})(v_{k}-v_{l})\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}\,dt.$
@@
${\mathbb{P}}(|X^{\prime}|\leq 1)\ll\int_{|t|\leq 1}\exp\Big\{-\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}dt.$ (33)
@@
$|{\mathcal{G}}|\geq{\varepsilon}n^{2},\qquad\sum_{(i,j)\in{\mathcal{G}}}(w_{i}-w_{j})^{2}\asymp_{{\varepsilon}}n.$ (34)
@@
$\sum_{k}D_{k}^{2}|{\mathcal{G}}_{k}|\asymp_{{\varepsilon}}n.$
@@
$\sum_{r\in{\mathcal{R}}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|^{2}_{{\mathbb{R}}/{\mathbb{Z}}}=\sum_{r\in{\mathcal{R}}}\Big|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big|^{2}\asymp_{\delta}\frac{t^{2}n^{4}D^{2}_{k}}{\Delta^{2}}.$

### v2

@@ 4.3. Diophantine Properties
Let $S_{\pi}((b_{ij}))=\sum_{i}b_{i\pi(i)}$, where $\pi$ is a uniform permutation. Then ${\mathbb{E}}S_{\pi}((b_{ij}))=(1/n)\sum_{i,j}b_{ij}$. It is also known [17, formula (89)] that
$\operatorname{Var}S_{\pi}((b_{ij}))=\frac{1}{4n^{2}(n-1)}\sum_{i\neq j;\,k\neq l}(b_{ik}-b_{jk}-b_{il}+b_{jl})^{2}.$ (18)
Hence, if $b_{ik}-b_{jk}-b_{il}+b_{jl}\in[-N,N]$, then $\operatorname{Var}(S_{\pi}((b_{ij})))\leq nN^{2}$. By Chebyshev’s inequality, with probability at least $1-16/C^{2}$,
$|S_{\pi}((b_{ij}))-{\mathbb{E}}S_{\pi}((b_{ij}))|\leq(C/4)\sqrt{n}\,N.$ (19)
For each fixed $1\leq s\leq r$, apply this bound to the sequence $b_{ij}=a_{ij;s}$ from (17). Taking the intersection over $s$, we obtain an event $\mathcal{E}$} (over the random permutation $\pi$) with ${\mathbb{P}}({\mathcal{E}})\geq 1-16r/C^{2}$ such that, simultaneously for all $s$,
$|S_{\pi}((a_{ij;s}))-\mu_{s}|\leq C\sqrt{n}\,N_{s},$
where $\mu_{s}={\mathbb{E}}S_{\pi}((a_{ij;s}))$.
Under this event $\mathcal{E}$}, the original sum $\sum_{i}a_{i\pi(i)}^{\prime}$ lies in the shifted GAP
$\big([\mu_{1}-C\sqrt{n}N_{1},\mu_{1}+C\sqrt{n}N_{1}]\cap{\mathbb{Z}}\big)g_{1}+\dots+\big([\mu_{r}-C\sqrt{n}N_{r},\mu_{r}+C\sqrt{n}N_{r}]\cap{\mathbb{Z}}\big)g_{r}.$
By the pigeonhole principle, there exists $x$ in this GAP such that
${\mathbb{P}}(S_{\pi}(a_{ij}^{\prime})=x)\geq\frac{{\mathbb{P}}({\mathcal{E}})}{(3C)^{r}n^{r/2}\prod_{i}N_{i}}\geq\rho$
for appropriate choices of the constants. ∎
@@
Embedding. The following theorem (see [60, Lemma 5.25], [42, Theorem 4.3]) allows us to assume that $a_{ij}$ are elements of ${\mathbb{F}}_{p}$ for some large prime $p$.
@@
Let $n^{{\varepsilon}}\leq n^{\prime}\leq n$. By averaging, at least a $(1-n^{\prime}/n)$-fraction of $a\in A$ (i.e. $(n^{4}-n^{\prime}n^{3})$ quadruples) satisfy
$\sum_{t\in L_{m}}\|\frac{at}{p}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\leq\frac{4m}{n^{\prime}}|L_{m}|.$
@@
$\sum_{t\in L_{m}}\|\frac{at}{p}\|^{2}_{{\mathbb{R}}/{\mathbb{Z}}}\leq l^{2}\frac{4m}{n^{\prime}}|L_{m}|.$ (20)
@@
Set $k=\sqrt{\frac{n^{\prime}}{160m}}$. By (20), we have $\bigcup_{l=1}^{k}lA^{\prime}\subset L_{m}^{*}$. Setting $A^{\prime\prime}=A^{\prime}\cup\{0\}$, this implies $kA^{\prime\prime}\subset L_{m}^{*}\cup\{0\}$. Hence
@@
The ambient field ${\mathbb{F}}_{p}$ is no longer important, so we may view the $a_{ij}$ as integers. We now invoke the following long-range inverse theorem (see [42, Theorem 3.2]).
@@
$k=\Omega\Big(\sqrt{\frac{n^{\prime}}{m}}\Big)=\Omega\Big(\sqrt{\frac{n^{\prime}}{\log n}}\Big),\quad\rho^{-1}\leq n^{C}\leq k^{2C/{\varepsilon}+1},$
(22) allows us to apply Theorem 5.3 with $\gamma=2C/{\varepsilon}+1$ and $X$ the set of distinct elements of $A^{{}^{\prime\prime}}$ (note that $kX=kA^{{}^{\prime\prime}}$ for $k\geq 2$). Hence $X$ is contained in a proper symmetric GAP $Q$ of rank $r=O_{C,{\varepsilon}}(1)$ and size
$O_{C,{\varepsilon}}(k^{-r}|kX|)=O_{C,{\varepsilon}}(k^{-r}|kA^{{}^{\prime\prime}}|)=O_{C,{\varepsilon}}\Big(\rho^{-1}e^{-m+2}(\sqrt{\frac{n^{\prime}}{m}})^{-r}\Big)=O_{C,{\varepsilon}}(\rho^{-1}(n^{\prime})^{-r/2}),$
@@
To complete the section we quickly deduce Theorem 3.1, Theorem 3.2 and Theorem 3.3.
@@
(of Theorem 3.1) Assume for contradiction that $\rho:=\sup_{x}{\mathbb{P}}\big(\sum_{i=1}^{n}w_{i}v_{\pi(i)}=x\big)\geq An^{-1/2}$ for some sufficiently large constant $A$. Set ${\varepsilon}=.1$ and $C=1/2$. By Theorem 2.1, there is a GAP $Q$ of rank $r$ and size $O(\rho^{-1}n^{-r/2})$ that contains a $(1-{\varepsilon})$-portion of the set of quadruples $\{(w_{i}-w_{j})(v_{k}-v_{l}):1\leq i,j,k,l\leq n\}$. This set must have at least $\Theta(1)$ (distinct) elements, because by assumption there are $\Theta_{c}(n^{2})$ pairs $i,j$ where $w_{i}-w_{j}\neq 0$ and $\Theta_{c}(n^{2})$ pairs $k,l$ where $v_{k}-w_{l}\neq 0$. Since $Q$ has rank at least $1$, we have $\Omega(n)=|Q|=O(\rho^{-1}n^{-1/2})=O(n/A)$, a contradiction if $A$ is sufficiently large. ∎
@@
(of Theorem 3.2) Assume for contradiction that $\rho:=\sup_{x}{\mathbb{P}}\big(\sum_{i=1}^{n}w_{i}v_{\pi(i)}=x\big)\geq An^{-3/2}$ for some sufficiently large constant $A$. Set ${\varepsilon}=.1$ and $C=3/2$. By Theorem 2.1, there is a GAP $Q$ of rank $r$ and size $O(\rho^{-1}n^{-r/2})$ that contains a $(1-{\varepsilon})$-portion of the set of quadruples $\{(w_{i}-w_{j})(v_{k}-v_{l}):1\leq i,j,k,l\leq n\}$. This set must have at least $\Theta(n)$ (distinct) elements, because by assumption there are $\Theta_{c}(n^{2})$ pairs $i,j$ where $w_{i}-w_{j}\neq 0$, and on average, each such pair $(i,j)$ is associated with $\Theta(n^{2})$ pairs $(k,l)$. Among these pairs, we just chose $\Theta(n)$ pairs of the form $(k_{0},l_{1}),\dots,(k_{0},l_{m}),m=\Theta(n)$. Since $Q$ has rank at least $1$, we have $\Omega(n)=|Q|=O(\rho^{-1}n^{-1/2})=O(n/A)$, a contradiction. ∎
Proof.
(of Theorem 3.3) Assume for contradiction that $\rho:=\sup_{x}{\mathbb{P}}\big(\sum_{i=1}^{n}w_{i}v_{\pi(i)}=x\big)\geq An^{-5/2}\log n$ for some sufficiently large constant $A$. Set ${\varepsilon}=.1$ and $C=5/2$. By Theorem 2.1, there is a GAP $Q$ of rank $r$ and size $O(\rho^{-1}n^{-r/2})$ that contains a $(1-{\varepsilon})$-portion of the set of quadruples $\{(w_{i}-w_{j})(v_{k}-v_{l}):1\leq i,j,k,l\leq n\}$. Denote this set by $S$. We claim that $S$ must have at least $\Theta(n^{2}/\log n)$ elements. Since $Q$ has rank at least $1$, this would imply $\Omega(n^{2}/\log n)=|Q|=O(\rho^{-1}n^{-1/2})=O(n^{2}/A\log n)$, a contradiction.
It remains to prove $|S|=\Omega(n^{2}/\log n)$. Let $A=\{v_{1},\ldots,v_{n},w_{1},\ldots,w_{n}\}$. For $x\in A-A$, let $r(x)$ denote the number of representations ab=xa-b=x with $(a,b)\in A^{2}$. By [5, Theorem 3], [^2] The statement in [5] is given for $A\subset{\mathbb{Z}}$, but the proof works verbatim for $A\subset{\mathbb{R}}$. we have
$\sum_{z}\Big|\sum_{xy=z}r(x)r(y)\Big|^{2}=\sum_{xy=x^{\prime}y^{\prime}}r(x)r(x^{\prime})r(y)r(y^{\prime})=O(|A|^{6}\log|A|).$
@@
$\frac{1}{|S|}\Big(\sum_{xy\in S}r(x)r(y)\Big)^{2}=O(|A|^{6}\log|A|).$
Moreover, $\sum_{xy\in S}r(x)r(y)$ counts at least the number of selected quadruples, namely $(1-{\varepsilon})n^{4}$. Therefore, $|S|=\Omega(n^{8}/|A|^{6}\log|A|)=\Omega(n^{2}/\log n)$, as claimed. ∎
6. Continuous setting: proof of Theorems 2.6 and 2.8
@@
Fact 6.1.
Let $I\subset[n]$ with $|I|\geq\delta n$ for some constant $\delta>0$. Then, there exists ${\mathcal{R}}\subset\{-n,\dots,n\}$ of size $|{\mathcal{R}}|=\Omega_{\delta}(n)$ such that for every $r\in{\mathcal{R}}$, there are $\Omega_{\delta}(n)$ pairs $x,y\in I$ with xy=rx-y=r.
Throughout this section, let $\mathcal{R}\subset\{-n,\dots,n\}$ be a set of size $|\mathcal{R}|=\Theta_{\delta}(n)$, as defined in Fact 6.1. Let $C$ and $A$ be positive constants, with $C$ sufficiently large relative to $\delta$ and $\varepsilon$, and $A$ sufficiently large relative to $C$.
@@
${\mathbb{P}}(|X|\leq 1)=O\Big(\Big|\int_{|t|\leq 1}k(t)\varphi_{X}(t)\,dt\Big|\Big)=O\Big(\int_{|t|\leq 1}|\varphi_{X}(t)|\,dt\Big),$ (23)
@@
${\mathbb{P}}(|X|\leq 1)=O\Big(\int_{|t|\leq 1}|\varphi_{S}(t)|\,dt\Big)=O\Big(\int_{|t|\leq 1}\exp\Big\{-\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}dt\Big).$ (24)
@@
$\sum_{r\in{\mathcal{R}}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\sum_{r\in{\mathcal{R}}}|t(w_{i}-w_{j})r|^{2}=\Theta_{\delta}(t^{2}n^{3}D_{k}^{2}).$
@@
$\sum_{r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}(n).$
@@
$\sum_{(i,j)\in{\mathcal{G}},r\in{\mathcal{R}}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega_{\delta,C}(1)\cdot\Big(\sum_{k:\,|t|D_{k}\leq 1/(Cn)}t^{2}n^{3}D_{k}^{2}|{\mathcal{G}}_{k}|+\sum_{k:\,|t|D_{k}>1/(Cn)}n|{\mathcal{G}}_{k}|\Big).$
@@
$\sum_{i,j,r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega_{\delta,C}(An^{2}\log n).$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}(t(w_{i}-w_{j})r)^{2}=\Theta_{\delta}(t^{2}n^{2}).$
@@
$\int_{|t|\leq(\sqrt{A\log n})/n}|\varphi_{S}(t)|\,dt\leq\int_{|t|\leq(\sqrt{A\log n})/n}\exp(-\Theta(t^{2}n^{2}))\,dt=O\Big(\frac{1}{n}\int_{{\mathbb{R}}}\exp(-\Theta(x^{2}))\,dx\Big)=O\Big(\frac{1}{n}\Big).$
@@
${\mathbb{P}}\Big(\Big|\sum_{i}w_{i}v_{\pi(i)}-L\Big|\leq\frac{1}{n}\Big)=O\left(\max\Big\{n^{-\sqrt{A}},\ \frac{1}{n}e^{-\Theta(L^{2})}\Big\}\right).$ (26)
@@
$\sum_{r\in{\mathcal{R}}}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{\delta}(n).$
@@
$\sum_{i,j,r}\|t(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega_{\delta}(n|{\mathcal{G}}|)=\Omega_{\delta}(A^{2}n^{2}\log n).$
@@
Our plan in the proof of (6) is to replace $k(t)$ in (23) by some smoother function that can be extended holomorphically to $\mathbb{C}$}. Our starting point is that
@@
$\int_{(\sqrt{A\log n})/n\leq|t|\leq\sqrt{A\log n}}e^{-\pi t^{2}}\varphi_{X}(t)dt\leq(2\sqrt{A\log n})n^{-2\sqrt{A}}=o(n^{-\sqrt{A}}).$
@@
$\Big|e^{-\pi(t-icL)^{2}/n^{2}}\Big|=e^{-\pi(t^{2}-c^{2}L^{2})/n^{2}}=\Theta(1),\quad\text{as }n\to\infty\text{ and }|t|\leq L=O(\sqrt{\log n}),$
@@
${\mathbb{P}}(|X|\leq 1)\leq\frac{e^{-cL^{2}}}{n}\cdot e^{C_{0}^{\prime}c^{2}L^{2}}\cdot 2L=O\Big(\frac{e^{-L^{2}/(16C_{0}^{\prime})}\,L}{n}\Big)=O\Big(\frac{e^{-\Theta(L^{2})}}{n}\Big).$
@@
We observe that the assumption $|w_{i}-w_{j}|\leq\frac{1}{A\sqrt{\log n}}$ in Theorem 2.6 cannot be relaxed to $|w_{i}-w_{j}|=O(\frac{1}{\sqrt{\log n}})$ when using only the characteristic function method (i.e., relying solely on Theorem 4.1). Indeed, suppose we partition $[n]$ into two disjoint sets $I\cup J=[n]$ with $|I|=2c\log n$ for some constant $c>0$. Let $w_{i}=\frac{1}{\sqrt{2c\log n}}$ and $\frac{1}{\sqrt{2c\log n}}$ for half of the $i\in I$ respectively, and $w_{j}=0$ for all $j\in J$. Then,
@@
$\sup_{x}{\mathbb{P}}(|S^{\prime}-x|\leq 1)=O\Big(\int_{|t|\leq 1}|\varphi_{S^{\prime}}(t)|\,dt\Big)=O\Big(\int_{|t|\leq 1}\exp\Big\{-\frac{1}{2n^{3}}\sum_{i,j,k,l}\Big\|\frac{tn^{3/2}}{\Delta}(w_{i}-w_{j})(v_{k}-v_{l})\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}\,dt\Big).$
@@
${\mathbb{P}}(|X^{\prime}|\leq 1)=O\Big(\int_{|t|\leq 1}\exp\Big\{-\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\Big\}dt\Big).$ (33)
@@
$|{\mathcal{G}}|\geq{\varepsilon}n^{2},\qquad\sum_{(i,j)\in{\mathcal{G}}}(w_{i}-w_{j})^{2}=\Theta_{{\varepsilon}}(n).$ (34)
@@
$\sum_{k}D_{k}^{2}|{\mathcal{G}}_{k}|=\Theta_{{\varepsilon}}(n).$
@@
$\sum_{r\in{\mathcal{R}}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|^{2}_{{\mathbb{R}}/{\mathbb{Z}}}=\sum_{r\in{\mathcal{R}}}\Big|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big|^{2}=\Theta_{\delta}\Big(\frac{t^{2}n^{4}D^{2}_{k}}{\Delta^{2}}\Big).$